# Associative Knowledge Graphs for Efficient Sequence Storage and Retrieval

### v1

@@ Abstract
This paper presents a novel approach for constructing associative knowledge graphs that are highly effective for storing and recognizing sequences. The graph is created by representing overlapping sequences of objects, as tightly connected clusters within the larger graph. Individual objects (represented as nodes) can be a part of multiple sequences or appear repeatedly within a single sequence. To retrieve sequences, we leverage context, providing a subset of objects that triggers an association with the complete sequence. The system’s memory capacity is determined by the size of the graph and the density of its connections. We have theoretically derived the relationships between the critical density of the graph and the memory capacity for storing sequences. The critical density is the point beyond which error-free sequence reconstruction becomes impossible. Furthermore, we have developed an efficient algorithm for ordering elements within a sequence. Through extensive experiments with various types of sequences, we have confirmed the validity of these relationships. This approach has potential applications in diverse fields, such as anomaly detection in financial transactions or predicting user behavior based on past actions.
Keywords structural associative knowledge graphs  $\cdot$ graph density  $\cdot$ sequence storage  $\cdot$ sequence retrieval  $\cdot$ context-based association
@@ 1 Introduction
Transformers [3], introduced later, eliminate the need for recurrent and convolutional neural networks, offering faster and more parallelizable structures. Hierarchical Temporal Memory (HTM), by [4] draws inspiration from the human brain’s structure. Projects like nupic.torch [5] integrate HTM principles into PyTorch highlighting the benefits of sparse representations.
Associative Long Short Term Memory LSTM [6] significantly improves d the performance of LSTM in tasks requiring storing and retrieving complex data by adding special associative memory. Clustered Neural Network (CNN) has the ability to recall stored messages quickly and enables full parallel processing, and its modification Restricted Clustered Neural Network (R-CNN) [7], [8] allows to reproduce data with fewer errors.
Knowledge Graphs offer a powerful framework for representing objects and their relationships, enabling complex decision-making [9]. Graph Neural Networks (GNNs [10]) has further enhanced the ability to process graph-structured data using deep learning techniques. The ANAKG (Active Neuro-Associative Knowledge Graphs [11], [12]) utilize dynamic neurons for learning sequences within an associative knowledge graph.
@@
A key challenge in applying structural approaches to sequence storage is the efficient ordering of retrieved sequence elements. Our key contribution lies in developing efficient algorithms for ordering retrieved sequence elements. This paper compares the efficiency of different sequence ordering algorithms and presents testing results on various datasets.
The program’s code is available on GitHub [14].
2 Building a Knowledge Graph
@@
In this work, we extensively analyze the structural properties of knowledge graphs, emphasizing the importance of knowledge graph density, measured by the ratio of used synaptic connections to the total possible synaptic connections among the nodes. We show that the theoretical analysis performed for scene memory [13] can be applied to sequence memory. Based on this analysis, we accurately estimate the most important dependencies between the number of nodes in the graph, graph density, context size, and the number of saved sequences. This allows us to determine and use the appropriate graph size or necessary context based on the size and number of sequences stored within the associative knowledge graph memory (referred to as graph memory).
A crucial aspect of reproducing a sequence is the proper ordering of its elements. An added challenge is that the given context does not necessarily preserve the order of the context elements in the stored sequence. A key contribution of this paper is demonstrating that the sparsity of the graph memory enables high memory capacity, which increases quadratically with the size of the graph.
Definition 1.
@@
Each transitive tournament will be placed in the memory graph (SSAKG). The individual elements of the sequence will be placed in the vertices of the graph with corresponding numbers. The order of the elements corresponds to the corresponding connections, i.e. the first element connects to all the others, the second to all the others except the first, and so on.
@@
As an example, let’s consider a SSAKG graph that contains 20 vertices and can store sequences consisting of numbers 1 to 20. Let’s assume that we first store the sequence [2,6,11] in this graph. The first element of this sequence (no. 2) is connected via the directed edges to the numbers 6 and 11. The second element (no. 6) connects to no. 11. No edge is derived from element no. 11 (Figure 1a). Figure 1b) shows the recorded sequence [2,6,11] and the sequence [11,8,2].
@@
When a knowledge graph becomes overly dense due to a limited number of unique objects, virtual objects can be a valuable solution. Virtual objects help maintain a sparse graph structure even when dealing with a large number of data. They are created by combining an object’s symbol with its location in the sequence. This limitation, the new approach presented in this work is necessary.
To further differentiate virtual objects, additional characteristics like color, position, or size can be incorporated. While it’s theoretically possible to reconstruct a sequence using the scene memory approach described in [13] if all subsequent elements are converted to virtual objects, this would require context elements to also include their precise sequence location. To avoid this limitation, the new approach presented in this work is necessary.
This section investigates the relationship between the density of a SSAKG graph and the quantity of stored sequences. As the number of sequences grows, so does the graph density influenced by factors such as average sequence length and the total number of nodes. The overlapping connections between densely connected clusters (known as transitive tournament subgraphs) contribute to the overall graph density. To address this overlap and maintain efficiency, we must consider the effective number of synaptic connections within a SSAKG that stores multiple transitive tournament subgraphs.
@@
where $l_{s}$ is the length of the added sequence.
Therefore, each new sequence written in the graph increases the graph density according to the relationship
@@
We can now easily connect the final density of the graph with the number of sequences $s$ stored in it. Since
$d_{0}=0$ (6)
we have the final graph density
@@
[…]
3 Creating a Structured Sequence Memory
@@
4 Sequence Elements Ordering Algorithms
To recover a sequence from an associative knowledge graph the retrieved elements must be ordered. This is particularly important for large datasets, as it allows us to convert the retrieved data into a properly ordered "transitive tournament graph matrix." While a simple sorting method might suffice for small, sparse graphs, more sophisticated techniques are needed for denser graphs containing elements from multiple sequences.
This section explores four algorithms for sequence element ordering:
@@ 4 Sequence Elements Ordering Algorithms
Node Ordering.
This algorithm aims to generate all possible orderings of rows and columns in matrix M to obtain an upper triangular matrix (without a diagonal) containing only ones in the upper right corner. It prioritizes rows with n-1 non-zero elements and removes both the identified row and its corresponding column. This process is repeated for all such rows. If rows have the same number of non-zero elements, all possible ordering combinations are considered. A key difference from Simple Sort is the removal of additional elements from potentially different sequences. For instance, when selecting the first row with n-1 non-zero elements, the entire corresponding column is removed, potentially containing elements from other sequences. Figure 2 illustrates the algorithm’s block diagram. To illustrate the results of the Node Ordering algorithm, let’s examine a straightforward example.
@@
Consider $55x$ 5 matrix M and assume that the path obtained by the Node Ordering algorithm consists of indices $[4,3,3,2,1]$. These indices represent the progressively eliminated rows (starting from row 4) of successively smaller matrices. Initially, with the original numbering $[1,2,3,4,5]$, we arrive at smaller and smaller matrices, which align with the original row numbers as illustrated here: $[1,2,3,5]\to[1,2,5]\to[1,2]\to[1]$. This path corresponds to the desired sequence elements renumbering: $[4,3,5,2,1]$.
It’s important to note that this procedure is reversible because, given the renumbered rows, we can determine the path generated by the Node Ordering algorithm. The block diagram of all the Node Ordering algorithms is shown in Fig.2 a), and Fig. 2 b) shows the structure of Prioritize functions that are different for each algorithm.
Figure 2: a) Block diagram of the Node Ordering. b) Prioritize functions that are different for each algorithm.
Enhanced Node Ordering.
@@
Weighted Edges Node Ordering.
@@
5 Comparison of Graph Ordering Algorithms Performance
Figure 3 compares the performance of all four algorithms in reconstructing sequences. The test involved 1000 sequences, each containing 15 elements. Figure 4a shows the results for a graph containing 1000 elements, while Figure 4b presents the results for a 2000-element graph.
In both cases, the context used for sequence reconstruction comprised 7 elements. Histograms illustrate the reproducibility of sequence elements. The x-axis represents the number of correctly reproduced elements, with 15 indicating perfect reconstruction. As evident, the Weighted Edges Ordering algorithm outperforms all others.
@@ 5 Comparison of Graph Ordering Algorithms Performance
Figure 4 explores the average number of node orderings obtained for sequences stored in SSAKG graphs of varying sizes (all sequences were 10 elements long) under the assumption that all sequence elements are known.
@@
As expected, the average number of orderings decreases with increasing graph size (represented on the lower x-axis) and increasing graph density (upper x-axis). This experiment focuses on Algorithms 2 (Node Ordering) and 4 (Weighted Edges Node Ordering) since Simple Sort performed significantly worse, and Enhanced Node Ordering offered similar results to Node Ordering. The figure clearly demonstrates that the Weighted Edges Node Ordering algorithm exhibits a much slower growth in the average number of node orderings, indicating its superior efficiency in accurately reconstructing sequences stored in the SSAKG graph.
Overall, these results highlight the importance of effective sequence node ordering algorithms for accurate data retrieval from associative knowledge graphs. The Weighted Edges Node Ordering algorithm demonstrates superior performance compared to simpler approaches.
6 Practical Examples
This section presents examples based on memorizing and recalling sentences from SSAKG memory. To obtain correct sentences, the program uses the NLTK library - Natural Language Toolkit. [15]. This library enables the reading of sentences and words with appropriate grammar for the specified language. Additionally, it contains a vast collection of texts grouped into “corpora". For our experiments, we utilized a corpus based on Gutenberg’s project [16], which includes several selected book entries. We carefully selected several thousand sentences with interesting parameters, filtering out common words (stop words) and punctuation marks [17].
@@ 6 Practical Examples
Sentences of Varying Lengths (10-15 words). We selected 1000 random sentences with lengths ranging from 10 to 15 words. The sentences were prepared using the previously described method. This random text contained 3037 different words. The word distributions in the text are depicted in Figure 6. The selected sentences are then encoded into SSAKG. Subsequently, using contexts of length 8,9, and 10, we attempted to recall these sentences from memory.
@@
Table 2: Correctness of Reading Sentences with 10-15 Words
@@
7 Conclusions

### v2

@@ Abstract
Background and Objective: The paper addresses challenges in storing and retrieving sequences in contexts like anomaly detection, behavior prediction, and genetic information analysis. Associative Knowledge Graphs (AKGs) offer a promising approach by leveraging sparse graph structures to encode sequences. The objective was to develop a method for sequence storage and retrieval using AKGs that maintain high memory capacity and context-based retrieval accuracy while introducing algorithms for efficient element ordering.
Methods: The study utilized Sequential Structural Associative Knowledge Graphs (SSAKGs). These graphs encode sequences as transitive tournaments with nodes representing objects and edges defining the order. Four ordering algorithms were developed and tested: Simple Sort, Node Ordering, Enhanced Node Ordering, and Weighted Edges Node Ordering. The evaluation was conducted on synthetic datasets consisting of random sequences of varying lengths and distributions, and real-world datasets, including sentence-based sequences from the NLTK library and miRNA sequences mapped symbolically with a window-based approach. Metrics such as precision, sensitivity, and specificity were employed to assess performance.
Results: The Weighted Edges Node Ordering algorithm demonstrated superior precision and resilience to graph density. In real-world applications, sentence retrieval achieved precision rates of 94.7%-97.3% for contexts of 8–10 words, while miRNA sequence retrieval using a 6-nucleotide window reached 99.6% precision at longer contexts. SSAKGs exhibited quadratic growth in memory capacity relative to graph size.
Conclusions: This study introduces a novel structural approach for sequence storage and retrieval. Key advantages include no training requirements, flexible context-based reconstruction, and high efficiency in sparse memory graphs. With broad applications in computational neuroscience and bioinformatics, the approach offers scalable solutions for sequence-based memory tasks.
Keywords sequence retrieval  $\cdot$ context-based association  $\cdot$ miRNA sequences  $\cdot$ associative knowledge graphs  $\cdot$ graph density  $\cdot$ ssakg package
@@ 1 Introduction
Transformers [3], introduced later, eliminate the need for recurrent and convolutional neural networks, offering faster and more parallelizable structures. Hierarchical Temporal Memory (HTM), by [4] draws inspiration from the human brain’s structure. Projects like NuPIC (nupic.torch) [5, 6, 7] integrate HTM principles into PyTorch highlighting the benefits of sparse representations.
Associative Long Short Term Memory LSTM [8] significantly improves the performance of LSTM in tasks requiring storing and retrieving complex data by adding special associative memory. Clustered Neural Network (CNN) can recall stored messages quickly and enables full parallel processing, and its modification Restricted Clustered Neural Network (R-CNN) [9, 10] allows to reproduce data with fewer errors.
Knowledge Graphs offer a powerful framework for representing objects and their relationships, enabling complex decision-making [11, 12]. Graph Neural Networks (GNNs [13]) has further enhanced the ability to process graph-structured data using deep learning techniques. The ANAKG (Active Neuro-Associative Knowledge Graphs [14, 15]) utilize dynamic neurons for learning sequences within an associative knowledge graph.
@@
  2. (ii)
    A key challenge in applying structural approaches to sequence storage is the correct ordering of retrieved sequence elements. Our key contribution lies in developing efficient algorithms for ordering retrieved sequence elements. This paper compares the efficiency of different sequence ordering algorithms and presents testing results on various datasets.
  3. (iii)
    The paper explicitly demonstrates its methods using biomedical datasets, such as miRNA sequences and linguistic datasets that mimic patient medical records.
  4. (iv)
    The SSAKG source code, Jupyter examples, and tutorials are available on GitHub [17]. The implemented software is also available as a ’ssakg’ package on the official Python Package Index (PyPI) repository (pypi.org/project/ssakg) [18].
2 Methods
2.1 Building a Knowledge Graph
@@
In this work, we extensively analyze the structural properties of knowledge graphs, emphasizing the importance of knowledge graph density, measured by the ratio of used synaptic connections to the total possible synaptic connections among the nodes. We show that the theoretical analysis performed for scene memory [16] can be applied to sequence memory. Based on this analysis, we accurately estimate the most important dependencies between the number of nodes in the graph, graph density, context size, and the number of saved sequences. This allows us to determine and use the appropriate graph size or required context based on the size and number of sequences stored within the associative knowledge graph memory (referred to as graph memory).
A crucial aspect of reproducing a sequence is the proper ordering of its elements. An additional challenge is that the given context does not necessarily preserve the order of the context elements in the stored sequence. A key contribution of this paper is demonstrating that the sparsity of the graph memory enables high memory capacity, which increases quadratically with the size of the graph.
Definition 1
@@
Each transitive tournament is embedded within the SSAKG memory graph. The individual elements of the sequence are placed as edges connecting the vertices of the graph, with edges labeled by corresponding numbers.
The order of the elements corresponds to the corresponding connections, i.e. the first element connects to all the others, the second to all the others except the first, and so on.
@@
As an example, let’s consider a SSAKG graph that contains 20 vertices and can store sequences consisting of numbers 1 to 20. Let’s assume that we first store the sequence [2,6,11] in this graph. The first element of this sequence (no. 2) is connected via the directed edges to the numbers 6 and 11. The second element (no. 6) connects to no. 11. No edge is derived from element no. 11 (Figure 1a). Figure 1b shows the recorded sequence [2,6,11] and the sequence [11,8,2].
@@
When a knowledge graph becomes overly dense due to a limited number of unique objects, virtual objects can be a valuable solution. Virtual objects help maintain a sparse graph structure even when dealing with a large number of data. They are created by combining an object’s symbol with its location in the sequence.
This section investigates the relationship between the density of a SSAKG graph and the quantity of stored sequences. Additional characteristics like color, position, or size can be incorporated to further differentiate virtual objects. While it’s theoretically possible to reconstruct a sequence using the scene memory approach described in [16] if all subsequent elements are converted to virtual objects, this would require context elements to also include their precise sequence location. To avoid this limitation, the new approach presented in this work is necessary.
As the number of sequences grows, so does the graph density influenced by factors such as average sequence length and the total number of nodes. The overlapping connections between densely connected clusters (known as transitive tournament subgraphs) contribute to the overall graph density. To address this overlap and maintain efficiency, we must consider the effective number of synaptic connections within an SSAKG that stores multiple transitive tournament subgraphs.
@@
where $l_{s}$ is the length of the added sequence. Therefore, each new sequence written in the graph increases the graph density according to the relationship
@@
We can now easily connect the final density of the graph with the number of sequences $s$ stored in it. Since $d_{0}=0$, we have the final graph density:
@@
In equation 7 we have the interdependence of the number of stored sequences and the graph density. This iterative formula converges quickly assuming density $d=0.5$.
2.2 Creating a Structured Sequence Memory
@@
2.3 Sequence Elements Ordering Algorithms
To recover a sequence from an associative knowledge graph the retrieved elements must be ordered. This is particularly important for large datasets, as it allows us to convert the retrieved data into a properly ordered "transitive tournament graph matrix." While a simple sorting method might suffice for small, sparse graphs, more sophisticated techniques are needed for denser graphs containing elements from multiple sequences. This section explores four algorithms for sequence element ordering:
@@ 2.3 Sequence Elements Ordering Algorithms
Node Ordering
This algorithm aims to generate all possible orderings of rows and columns in matrix M to obtain an upper triangular matrix (without a diagonal) containing only ones in the upper right corner. It prioritizes rows with n-1 non-zero elements and removes both the identified row and its corresponding column. This process is repeated for all such rows. If rows have the same number of non-zero elements, all possible ordering combinations are considered. A key difference from Simple Sort is the removal of additional elements from potentially different sequences. For instance, when selecting the first row with n-1 non-zero elements, the entire corresponding column is removed, potentially containing elements from other sequences. The pseudocode of algorithm is shown in Algorithm 1. To illustrate the results of the Node Ordering algorithm, let’s examine a straightforward example.
@@
Consider $5x5$ matrix M and assume that the path obtained by the Node Ordering algorithm consists of indices $[4,3,3,2,1]$. These indices represent the progressively eliminated rows (starting from row 4) of successively smaller matrices. Initially, with the original numbering $[1,2,3,4,5]$, we arrive at smaller and smaller matrices, which align with the original row numbers as illustrated here: $1,2,3,5]\to[1,2,5]\to[1,2]\to[1]$. This path corresponds to the desired sequence elements renumbering: $[4,3,5,2,1]$.
It’s important to note that this procedure is reversible because, given the renumbered rows, we can determine the path generated by the Node Ordering algorithm. The pseudocode of the Node Ordering is shown in Algorithm 1, Algorithm 2 shows the structure of Prioritize functions that are different for each algorithm.
node_ordering (M,path,set_of_paths(M,path,set\_of\_paths)
 Input: M - unordered adjacency matrix of sequence
 Input: path - current path
 Input: set_of_paths - set of founded paths
 Output: set of founded paths
   n $\leftarrow$ dimension(M)
 if $n>0$ then
    // In ideal case there is one row
      non_zeros_ $\leftarrow$ find rows contain (n1)(n-1) non-zeros elements
    // Prioritize function is different for each algorithm
      prioritized_indices $\leftarrow$ prioritize(M,non_zeros_indices)
    foreach i in prioritized_indices do
         path.append(i)
       // The algorithm has found the prioritized rows in the matrix. We can look for the next ones.
         M $\leftarrow$ remove row i and column i from matrix M.
       node_ordering(M,path,set_of_paths(M,path,set\_of\_paths)
 else
      set_of_paths.append(path)
    return set_of_paths
Algorithm 1 Pseudocode of the Node Ordering.
Enhanced Node Ordering
@@
Weighted Edges Node Ordering
@@
// Inputs and outputs for all prioritize functions
Input: M - unordered adjacency matrix of sequence
Input: non_zeros_indices - set of selected indices
Output: prioritized indices
// Node ordering - empty function
prioritize (M,non_zeros_indices(M,non\_zeros\_indices)
   prioritized_indices $\leftarrow$ non_zeros_indices
 return prioritized_indicesprioritized\_indices
// Enchanced node ordering
prioritize (M,non_zeros_indices)(M,non\_zeros\_indices)
   Calculate weights: $w_{ij}=\frac{1}{M_{ij}}$
   Summarize weights in each row
   prioritized_indices $\leftarrow$ find rows with min weights
 return prioritized_indices
// Weighted edges node ordering
prioritize (M,non_zeros_indices)(M,non\_zeros\_indices)
 // Determine the prioritised number, in this case the number is equal to the dimension of M.
   prioritized_number $\leftarrow$ dimension(M)
   prioritized_indices $\leftarrow$ find rows with majority of prioritized numbers
 return prioritized_indices
Algorithm 2 Prioritize functions.
3 Results
3.1 Comparison of Graph Ordering Algorithms Performance
The following example uses randomly generated data as test sequences. When generating the example sequences, we used a uniform distribution. Figure 2 compares the performance of all four algorithms in reconstructing sequences. The test involved 1000 sequences, each containing 15 elements. Figure 4a shows the results for a graph containing 1000 elements, while Figure 4b presents the results for a 2000-element graph.
The tests were performed on an Apple Mac Mini M1 with 16GB RAM. Depending on the complexity of the test, the duration of a single test ranged from 1 second to 1 minute.
In both cases, the context used to reconstruct the sequence included 7 unordered elements. Histograms illustrate the reproducibility of sequence elements. The x-axis represents the number of correctly reproduced elements, with 15 indicating perfect reconstruction. As evident, the Weighted Edges Ordering algorithm outperforms all others.
@@ 3.1 Comparison of Graph Ordering Algorithms Performance
Figure 3 explores the average number of node orderings obtained for sequences stored in SSAKG graphs of varying sizes (all sequences were 10 elements long) assuming that all sequence elements are known.
@@
As expected, the average number of orderings decreases with increasing graph size (represented on the lower x-axis) and increasing graph density (upper x-axis). This experiment focuses on Algorithms 2 (Node Ordering) and 4 (Weighted Edges Node Ordering) since Simple Sort performed significantly worse, and Enhanced Node Ordering offered similar results to Node Ordering. The figure demonstrates that the Weighted Edges Node Ordering algorithm exhibits a much slower growth in the average number of node orderings, indicating its superior efficiency in accurately reconstructing sequences stored in the SSAKG graph.
Overall, these results highlight the importance of effective sequence node ordering algorithms for accurate data retrieval from associative knowledge graphs. The Weighted Edges Node Ordering algorithm demonstrates superior performance compared to simpler approaches. This is essential for tasks like ordering genetic sequences (e.g., miRNA) or reconstructing medical event timelines.
Computational complexity. The space complexity exhibits an asymptotic growth rate of $O(n^{2})$ (where n is the number of all available symbols) and is independent of the length of the sequences (l parameter). This is because the complement matrix’s space requirement depends solely on the number of symbols (n) and remains constant, irrespective of the sequence length (l) or the number of stored sequences. Regarding time complexity, the asymptotic growth is $O(n)$. While the dependence on the sequence length is challenging to estimate precisely, it is approximately $O(l^{2})$.
The number of stored sequences $n$ does not affect the read speed visibly. However, the main issue is the length of the sequence $l$. The density $d$ of the graph increases proportionally to the square of the length of the sequence $l$.
Regarding computational requirements, SSAKG can handle numbers of over 100,000 symbols (tested on an NVIDIA DGX A100 computational server).
3.2 Practical Examples: Encoding Sentences in SSAKG
This section presents examples based on memorizing and recalling sentences from SSAKG memory. The program uses the NLTK library - Natural Language Toolkit to obtain correct sentences. [19]. This library enables reading sentences and words with appropriate grammar for the specified language. Additionally, it contains a vast collection of texts grouped into "corpora". For our experiments, we utilized a corpus based on Gutenberg’s project [20], which includes several selected book entries. We randomly selected several thousand sentences with interesting parameters, filtering out common words (stop words) and punctuation marks [21]. To tokenise and remove stop words, we used the appropriate tools provided by the NLTK library.
@@ 3.2 Practical Examples: Encoding Sentences in SSAKG
Sentences of Varying Lengths (10-15 words). We selected 1000 random sentences with lengths ranging from 10 to 15 words. The sentences were prepared using the previously described method. This random text contained 3037 different words. The word distributions in the text are depicted in Figure 5. The selected sentences are then encoded into SSAKG. Subsequently, using contexts of lengths 8,9, and 10, we attempted to recall these sentences from memory.
@@
Table 2: Correctness of Reading Sentences with 10–15 Words
@@
3.3 Practical Examples: Encoding MicroRNA Sequences in SSAKG
MicroRNAs (miRNAs) are small, non-coding RNAs that regulate gene expression by binding to target mRNA molecules. While computational methods can predict miRNA-mRNA interactions, inconsistencies between databases remain a challenge. Although multiple online databases exist, they often report inconsistent results regarding miRNA targets, as target sequences can differ between databases [22, 23]. Therefore, new and improved methods for matching miRNAs to mRNA sequences are continuously being developed. Our approach efficiently identifies miRNA sequences, focusing on their unique characteristics. This approach can also be adapted for matching miRNAs with mRNA targets.
Example 5.
We selected a dataset of 1000 random miRNA sequences each 20 nucleotide long. The MicroRNA data was downloaded from miRbase [24]. The challenge lies in representing these sequences efficiently in an SSAKG. A straightforward approach would be to represent each nucleotide as a node in the graph. However, this would lead to a limited number of unique symbols (four, corresponding to the four nucleotide bases: A, C, G, T). To increase the diversity of symbols and capture more complex sequence patterns, we introduced a window-based approach. Symbols are created based on a given nucleotide and the nucleotides surrounding it (Figure 6). Each nucleotide is numbered as shown in Figure 6 (a). The sequence of numbers within the window shown in Figure 6 (b) is then converted into a single symbol using a quadruplex number system Figure 6 (c). This process effectively multiplies the number of possible symbols and allows for more nuanced representation of sequence information in the SSAKG.
Figure 6: The mapping algorithm for converting nucleotide sequences into symbol sequences. (a) Nucleotide symbols. (b) Construction of corresponding window. (c) Calculating a symbol from a quadruplex system
The results are presented in Table 3. A window containing 6 nucleotides was selected in the test. By using this approach, we were able to significantly increase the diversity of symbols in the SSAKG, enabling more efficient storage and retrieval of sequence information. For miRNA test we created a context in with consecutive elements, as opposed to the standard context, which contains unsorted elements. Increasing the number of sequences in a dataset leads to a decrease in the accuracy of sequence reconstruction from context, which is due to the lack of uniqueness of context in this larger dataset.
Table 3: Correctness of Reading miRNA sequences with 20 Nucleotides
Context Length Correct Sets of Nucleotides Correct Order
3 64.10% 94.54%
4 79.40% 95.09%
5 86.60% 94.80%
6 91.40% 94.97%
9 95.70% 94.67%
Example 6.
We selected a dataset containing more than 8,000 miRNA sequences of 21 nucleotides each. These are all sequences of this length currently available on miRBase. A window containing 7 nucleotides was selected in the test. The results are presented in Table 4.
Table 4: Correctness of Reading miRNA sequences with 21 Nucleotides
Context Length Correct Sets of Nucleotides Correct Order
5 63.04% 82.25%
6 74.30% 82.37%
9 84.54% 81.73%
The paper introduces a window-based symbolic representation for miRNA sequences, increasing the diversity of symbols used in the graph. This method significantly improves storage and retrieval efficiency for nucleotide sequences, crucial for genomic studies and disease research. Precision values for retrieving miRNA sequences reached 95.7% with a 9-context length.
4 Discussion and conclusion